OpenStudy (anonymous):

you are studying for your final exam of the semester. up to this point,you recieved 3 exam scores of 83%,85%, and 84%. to recieve a grade of B in the class you must average exam score between 80% and 89% for all 4 exams including the final. Find the widest range of scores that you can get on the final exam in order to recieve a grade of B for the class.Use interval notation

OpenStudy (anonymous):

To find the mean/average score, you need to add all of the scores and divide by the number of scores you have. If you make the final exam (the unknown), s, then you get the following formula for the final average: (83+85+84+s)/4=f, where f=final score Simplified (add the already known scores together), you get, (252+s)/4=f The lowest final score is 80 and the highest is 89, so you just have to find the value of s when f=80 and f=89 as follows: (252+s)/4=80 and (252+2)/4=89 Need more, just ask.

OpenStudy (anonymous):

i can only choose [68,100] [65,97] [62,100]

OpenStudy (anonymous):

Are you able to find the value of s for the two equations? I just noticed a mistake in the second one, should be (252+s)/4=89 I'm not sure how that 2 got in there.

OpenStudy (anonymous):

i tried using [68,100]

OpenStudy (anonymous):

nvm i got it thanks tho